# Differential fuzz testing to detect tampering in sensor systems and its application to arms control authentication

### v1

@@ 2.4 Attack pathways
In this work, we consider two relatively simple attack mechanisms, both based on the same point on the attack surface: an attack executed on the listmode gamma ray data stream by overwriting core functions of a loaded numpy Python library that will be called later at runtime. This open-source library is a general-purpose package for manipulating multidimensional arrays, and widely used for scientific computing. The radiation data acquisition code uses routines in numpy to perform array copy (numpy.copy) and disk output operations (numpy.save). To substitute these routines with the attacker’s implementation, we create and hide a wrapper script that will be implicitly called with the Python interpreter and update the original function behavior at runtime. This technique is known as monkey patching and is similar to the Stuxnet computer worm attack, which replaced a dynamic library (DLL) used in Siemens Step 7 software for controlling centrifuges with a malicious copy [37, 38]. Our attack on the instrument data stream can be executed without modifying the analysis or data acquisition code itself, which would be easily detectable by source code hashing. More sophisticated attacks are discussed in Section 4.
@@ 2.4.1 Time-based attack
Fig. 2 shows an example scenario in which the inspector and host have agreed to make warhead verification measurements on a given Friday between 22223333 pm. Before sending the inspection system to the host country, the inspector runs an initial fuzz sequence to establish the baseline behavior of the system; one of the fuzzed parameters is the system time. During certification at the host facility, the host monkey patches the analysis library in order to randomly duplicate listmode gamma ray events with a $10\%$ probability. To avoid detection by simple source code hashing, the host ensures that this malicious injected script is well-hidden on the file system and that the redirected malicious routine(s) will modify the data only if the system time is between 22223333 pm on Friday. The host then prepares a hoax weapon with some of its nuclear material removed, such that its gamma emission rate is reduced by $10\%$, and retains the removed nuclear material in secret. Then, during the designated Friday 22223333 pm window, the inspector and host make treaty verification measurements. The attack triggers, and the inspector is fooled into believing the correct amount of weapons material has been accounted for. However, after the treaty verification measurements, the inspector runs another fuzz sequence with the same input parameters as the baseline. Because the fuzzed parameters include the system time, the malicious code is occasionally re-executed without the corresponding reduction in source material, producing an observable change in the output gamma spectra that is detected by the $\chi^{2}/\nu$ metric.
@@ 2.4.2 Counts-based attack
Once again, this attack can be detected by the fuzz testing procedure. Fuzzing detector parameters such as the high voltage, fine gain, and pulse width can change the gross counts in the spectrum through various mechanisms such as shifting higher-energy parts of the spectrum outside the dynamic range of the ADC or increasing the number of low-energy counts recorded (regardless of whether the counts are real or for example electronic noise). During the second fuzz testing sequence, the malicious logic detects off-normal gross count values due to the off-normal parameter values, rather than the presence of a hoax warhead. The attack executes, modifying the listmode counts to compensate for the difference in total counts, creating large differences in the $\chi^{2}/\nu$ metric between the first and second fuzzing output sequences.
@@ 3.1 Time-based attack
Figs. 4 and 5 show how our fuzz testing procedure detects the time-based attack on our demonstration NaI gamma spectroscopy system. We first establish the baseline system behavior via $100$ separate radiation measurements with a constant dwell time of $30$ s and PMT high voltage of $1000$ V, while fuzzing the pulse width uniformly (pseudo-)randomly between $0.75$ and $2.0$ µs, the fine gain between $0.5$ and $1.2$, and the system time between $1$ pm and $4$ pm. The left panel of Fig. 4 shows the $\chi^{2}/\nu$ metric as a function of measurement number when the pseudorandom fuzz random inputs are repeated on a system that has not been tampered with; all $\chi^{2}/\nu$ values are below the threshold of $2$, and no alarm is raised. Conversely, the right panel shows the system behavior when the fuzz random inputs are repeated on the tampered system. Approximately one third of the fuzzed system times fall within the 22223333 pm attack window and trigger the malicious code that artificially duplicates counts with a $10\%$ probability. As shown in Fig. 5, this leads to a corresponding $10\%$ increase in counts across the entire energy spectrum, and a $\chi^{2}/\nu>2$, raising an alarm.
@@ 4 Discussion
First, more sophisticated attacks may expand the state space of the system and will require more effort to detect. For instance, the attacker can code their malicious program to only activate under certain specific conditions, also known as a logic bomb—historical analogues here include Stuxnet [37, 38], the Volkswagen diesel emissions scandal [39], and the arms-control-focused 2015 Underhanded C Contest [40]. Our time-based attack is a straightforward case in which an activation condition (a given system time range) is present, but is also one of the parameters that is directly fuzzed, and thus the attack is readily detectable. In our counts-based attack, the attack uses a small amount of “live” data (namely the count rate) to determine whether to activate the attack, but is still detected because such conditions are replicated during the fuzz testing. A more advanced attack might activate only when the system determines via gamma ray spectroscopic analysis that the system is measuring special nuclear material (i.e., weapons-grade U and/or Pu). Then, if the fuzz testing measurements only use (say) Cs-137 (as they do in this work), the attack will never be triggered during the fuzz testing measurements and will not be detected by the fuzz testing analysis. It may become necessary, then, to perform the fuzz testing using standardized U/Pu calibration sources, though this will increase the operational challenge compared to a Cs-137 check source.
@@
Systematic environmental changes between different fuzz sequences present an additional challenge, as in this work we have either (1) assumed that variations in output gamma spectra can only arise from Poisson noise or tampering, or (2) adjusted $\chi^{2}/\nu$ thresholds in a somewhat ad-hoc fashion to account for environmental variations. Our NaI+PMT detection system is particularly sensitive to environmental variations due to the non-linearity of the PMT gain with temperature. Such variations stretch or squeeze the gamma spectra in channel space, potentially causing spectral misalignments and thus anomalously large $\chi^{2}/\nu$ values. Similarly, variable natural backgrounds across time and measurement locations may also increase spectral disagreements, but the influence of background can be minimized with shielding and strong calibration sources. In this work, we have tried to reduce the impact of environmental variations by collecting both fuzz sequences in the same location within a few hours, and by setting a large enough alarm threshold such that the environmental variation over this time span does not cross the threshold. We are also exploring gain stabilization methods, but we note that these add complexity to the system and in fact expand the attack surface. In future arms control treaties, therefore, detector materials with better environmental robustness (such as CdZnTe [41]) may be preferred for fuzz testing applications.
@@
We emphasize again that our demonstration system is only a proxy for a real arms control verification system, as it is unlikely that a general-purpose personal computer running a full operating system and suite of other applications with a multitude of third-party libraries can pass the authentication and certification requirements for in-field use around nuclear warheads. While current inspection equipment typically consists of (relatively) simple analogue electronic counters designed for a specific measurement, more advanced verification methods could require (say) a small gamma spectroscopy software library running on a microcontroller or edge computer. If software is present, the source code, binaries, and compiler should be available to the inspector for (necessary but not sufficient) authentication checks against both straightforward modifications and advanced attacks such as Trusting Trust [42] or the xz hack [43].
@@
More generally, further work will be necessary to determine the role of fuzz testing within the overall authentication CONOPS. In ongoing work, for instance, we intend to integrate our radiation detection system with the Chain of Custody Item Monitor (CoCIM) [44], a fiber-optic active seal developed for use in warhead verification applications, and deploy fuzz testing on the CoCIM software. Conversely, the utility of fuzz testing for information barriers (e.g., [45, 46, 47, 48, 49]) may be limited, since they typically reduce sensitive measurement data to 11116666 output bits of non-sensitive information. We also note that many CONOPS details will likely emphasize ease-of-use for the inspector and not necessarily just the detection power of the fuzz testing approach. As a simple example, the inspector may not want (or may not be able) to run fuzz testing over millions of parameter combinations over the span of many weeks, thereby limiting the possible coverage of the fuzz testing approach. In addition, warhead verification systems such as TRIS [35] are typically designed to have little or no easily-manipulable user inputs as a security feature, which may make fuzz testing operationally challenging. It is important to emphasize therefore that fuzz testing will not make cheating on an arms control treaty impossible, especially not on its own. Rather, it should be deployed as one among many tools for making cheating infeasibly difficult and/or expensive.

### v2

@@ 2.4 Attack pathways
In this work, we consider two relatively simple attack mechanisms, both based on the same point on the attack surface: an attack executed on the listmode gamma ray data stream by overwriting core functions of a loaded numpy Python library that will be called later at runtime. This open-source library is a general-purpose package for manipulating multidimensional arrays, and widely used for scientific computing. The radiation data acquisition code uses routines in numpy to perform array copy (numpy.copy) and disk output operations (numpy.save). To substitute these routines with the attacker’s implementation, we create and hide a wrapper script that will be implicitly called with the Python interpreter and update the original function behavior at runtime. This technique is known as monkey patching and is similar to the Stuxnet computer worm attack, which replaced a dynamic library (DLL) used in Siemens Step 7 software for controlling centrifuges with a malicious copy [langner2013kill, zetter2014stuxnet]. Our attack on the instrument data stream can be executed without modifying the analysis or data acquisition code itself, which would be easily detectable by source code hashing. In particular, the attacks drop or duplicate listmode gamma ray event data, depending on the attacker’s goal. For example, the attacker may submit an item with some of its gamma-emitting material removed, and duplicate listmode counts to compensate; in this work, we simulate such a reduction of nuclear material by simply increasing the source-to-detector distance. More sophisticated attacks are discussed in Section 4.
@@ 2.4.1 Time-based attack
Fig. 2 shows an example scenario in which the inspector and host have agreed to make warhead verification measurements on a given Friday between 2233 pm. Before sending the inspection system to the host country, the inspector runs an initial fuzz sequence to establish the baseline behavior of the system; one of the fuzzed parameters is the system time. During certification at the host facility, the host monkey patches the analysis library in order to randomly duplicate listmode gamma ray events with a $10\%$ probability. To avoid detection by simple source code hashing, the host ensures that this malicious injected script is well-hidden on the file system and that the redirected malicious routine(s) will modify the data only if the system time is between 2233 pm on Friday. The host then prepares a hoax weapon with some of its nuclear material removed, such that its gamma emission rate is reduced by $10\%$, and retains the removed nuclear material in secret. Then, during the designated Friday 2233 pm window, the inspector and host make treaty verification measurements. The attack triggers, and the inspector is fooled into believing the correct amount of weapons material has been accounted for. However, after the treaty verification measurements, the inspector runs another fuzz sequence with the same input parameters as the baseline. Because the fuzzed parameters include the system time, the malicious code is occasionally re-executed without the corresponding reduction in source material, producing an observable change in the output gamma spectra that is detected by the $\chi^{2}/\nu$ metric.
@@ 2.4.2 Counts-based attack
Once again, this attack can be detected by the fuzz testing procedure. Fuzzing detector parameters such as the high voltage, fine gain, and pulse width can change the gross counts in the spectrum through various mechanisms such as shifting higher-energy parts of the spectrum outside the dynamic range of the ADC or increasing the number of low-energy counts recorded (regardless of whether the counts are real or are for example electronic noise). During the second fuzz testing sequence, the malicious logic detects off-normal gross count values due to the off-normal parameter values, rather than the presence of a hoax warhead. The attack executes, modifying the listmode counts to compensate for the difference in total counts, creating large differences in the $\chi^{2}/\nu$ metric between the first and second fuzzing output sequences.
@@ 3.1 Time-based attack
Figs. 4 and 5 show how our fuzz testing procedure detects the time-based attack on our demonstration NaI gamma spectroscopy system. We first establish the baseline system behavior via $100$ separate radiation measurements with a constant dwell time of $30$ s and PMT high voltage of $1000$ V, while fuzzing the pulse width uniformly (pseudo-)randomly between $0.75$ and $2.0$ µs, the fine gain between $0.5$ and $1.2$, and the system time between $1$ pm and $4$ pm. The left panel of Fig. 4 shows the $\chi^{2}/\nu$ metric as a function of measurement number when the pseudorandom fuzz random inputs are repeated on a system that has not been tampered with; all $\chi^{2}/\nu$ values are below the threshold of $2$, and no alarm is raised. Conversely, the right panel shows the system behavior when the fuzz random inputs are repeated on the tampered system. Approximately one third of the fuzzed system times fall within the 2233 pm attack window and trigger the malicious code that artificially duplicates counts with a $10\%$ probability. As shown in Fig. 5, this leads to a corresponding $10\%$ increase in counts across the entire energy spectrum, and a $\chi^{2}/\nu>2$, raising an alarm. In this simple example, the chosen threshold of $\chi^{2}/\nu=2$ results in a false alarm rate of 0 and a false negative rate of $1\%$.
@@ 4 Discussion
First, more sophisticated attacks may expand the state space of the system and will require more effort to detect. For instance, the attacker can code their malicious program to only activate under certain specific conditions, also known as a logic bomb—historical analogues here include Stuxnet [langner2013kill, zetter2014stuxnet], the Volkswagen diesel emissions scandal [schiermeier2015science], and the arms-control-focused 2015 Underhanded C Contest [underhanded_c_contest]. Our time-based attack is a straightforward case in which an activation condition (a given system time range) is present, but is also one of the parameters that is directly fuzzed, and thus the attack is readily detectable. In our counts-based attack, the attack uses a small amount of “live” data (namely the count rate) to determine whether to activate the attack, but is still detected because such conditions are replicated during the fuzz testing. A more advanced attack might activate only when the system determines via gamma ray spectroscopic analysis that the system is measuring special nuclear material (i.e., weapons-grade U and/or Pu). Then, if the fuzz testing measurements only use (say) Cs-137 (as they do in this work), the attack will never be triggered during the fuzz testing measurements and will not be detected by the fuzz testing analysis. It may become necessary, then, to perform the fuzz testing using standardized U/Pu calibration sources, though this will increase the operational challenge compared to a Cs-137 check source. Similarly, the attack code could be designed to determine whether the system is in fuzz testing mode or weapon verification mode based on the measurement dwell time(s)—the repeated constant $30$ s dwell time used here for simplicity could indicate fuzz testing while longer dwells might indicate weapon verification. Fuzzing the measurement dwell times, as suggested in Table 1, may provide an additional layer of “anti-anti-tamper” defense.
A malicious host may also attempt a replay attack, in which they construct an approximate oracle that provides realistic-looking but synthetic spectra at a wide range of fuzz input parameters, bypassing the detector+readout chain entirely. Such an attack could be based on a digital twin of the detector system, or may comprise a data-driven interpolation table or other latent space representation. Both methods will require additional compute resources that may be difficult to hide from the inspector, and will suffer from some level of approximation error that would be interesting to quantify in future work. Moreover, even a well-crafted and well-hidden replay attack could be discovered by the addition of fuzz input parameters that cannot be controlled by the detection system, and thus cannot be modeled by the attack. For instance, fuzzing the source-to-detector distance could detect a replay attack that assumes the calibration source is located at a fixed distance of $d=1$ m, since the incident signal intensity varies as $1/d^{2}$.
@@
Systematic environmental changes between different fuzz sequences present an additional challenge, as in this work we have either (1) assumed that variations in output gamma spectra can only arise from Poisson noise or tampering, or (2) adjusted $\chi^{2}/\nu$ thresholds in a somewhat ad-hoc fashion to account for environmental variations. Our NaI+PMT detection system is particularly sensitive to environmental variations due to the non-linearity of the PMT gain with temperature. Such variations stretch or squeeze the gamma spectra in channel space, potentially causing spectroscopic misalignments and thus anomalously large $\chi^{2}/\nu$ values. Similarly, variable natural backgrounds across time and measurement locations may also increase spectroscopic disagreements, but the influence of background can be minimized with shielding and strong calibration sources. In this work, we have tried to reduce the impact of environmental variations by collecting both fuzz sequences in the same location within a few hours, and by setting a large enough alarm threshold such that the environmental variations over this time span do not cross the threshold. One could also explore gain stabilization methods, but we note that these add complexity to the system and in fact expand the attack surface. In future arms control treaties, therefore, detector materials with better environmental robustness (such as CdZnTe [park2010effect]) may be preferred for fuzz testing applications. More generally, it is important to reduce these kinds of systematic variations in any verification method so as to maintain a strict alarm threshold and thus reduce the opportunity for an attacker to hide an attack within the observed noise.
@@
We emphasize again that our demonstration system is only a proxy for a real arms control verification system, as it is unlikely that a general-purpose personal computer running a full operating system and suite of other applications with a multitude of third-party libraries can pass the authentication and certification requirements for in-field use around nuclear warheads. While current inspection equipment typically consists of (relatively) simple analogue electronic counters designed for a specific measurement, more advanced verification methods could require (say) a small gamma spectroscopy software library running on a microcontroller or edge computer. If software is present, the source code, binaries, and compiler should be available to the inspector for (necessary but not sufficient) authentication checks against both straightforward modifications and advanced attacks such as Trusting Trust [thompson1984reflections] or the xz hack [cox2024xz].
@@
Testing may also be accelerated through the use of so-called “gray-box” fuzz testing techniques, such as power scheduling and coverage-based feedback design. However, in the arms control application space especially, we expect that the far-simpler black-box technique (uniformly randomly sampling the fuzz parameter space) will be easiest for treaty partners to agree on.
More generally, further work will be necessary to determine the role of fuzz testing within the overall authentication CONOPS. In ongoing work, for instance, we intend to integrate our radiation detection system with the Chain of Custody Item Monitor (CoCIM) [brotz2017non], a fiber-optic active seal developed for use in warhead verification applications, and deploy fuzz testing on the CoCIM software. Conversely, the utility of fuzz testing for information barriers (e.g., [close2001information, chambers2010uk, yan2015nuclear, kutt2018information, brotz2023development]) may be limited, since they typically reduce sensitive measurement data to 1166 output bits of non-sensitive information. We also note that many CONOPS details will likely emphasize ease-of-use for the inspector and not necessarily just the detection power of the fuzz testing approach. As a simple example, the inspector may not want (or may not be able) to run fuzz testing over millions of parameter combinations over the span of many weeks, thereby limiting the possible coverage of the fuzz testing approach. In addition, expanding the technique to different systems presents operational challenges. Although the basic fuzzing concept remains constant across systems, for each new system one must identify a set of fuzzable input parameters that covers a significant portion of the system’s state space. For example, while nearly all radiation detection equipment will have some kind of high voltage and pulse width parameters (Table 1), warhead verification systems such as TRIS [seager2001trusted] are typically designed to have few or no easily-manipulable user inputs as a security feature, which may make such parameter identification challenging. Finally, as alluded to in Section 2.1, fuzz testing does not address, and in some cases may even exacerbate the certification problem, at the expense of addressing the authentication problem. It is important to emphasize therefore that fuzz testing will not make cheating on an arms control treaty impossible, especially not on its own. Rather, it should be deployed as one among many tools for making cheating infeasibly difficult and/or expensive.